OpenStudy (anonymous):

The rate of effusion of oxygen to an unknown gas is 0.935.What is the other gas?

OpenStudy (aaronq):

us graham's law of effusion: \[\frac{ v _{O _{2}} }{v _{unknown}}=\sqrt{\frac{ M _{unknown} }{ M _{O _{2}} }}\]